OpenStudy (anonymous):

A 50 kg person is standing on a bathroom scale in an elevator moving upward. a.) if the scale reads 56 kg, what is the magnitude of upward acceleration of the elevator (m/s^2)? b.) Now the elevator is traveling up while slowing down. If the magnitude of deceleration is the same as that in part (a), what is reading given by the scale?

OpenStudy (anonymous):

Use the fact that the scale is "programmed" by its construction to measure the force of gravity (in Newtons) and automatically divide by g, presenting the gravitational mass in kg on its face.

OpenStudy (anonymous):

What did the R stand for and how did you get it to be 560?

OpenStudy (anonymous):

accelerating up: Normal= mg +ma Normal=50g +50a =560N 50a =60 a=(6/5) m/s^2 accelerating down: Normal= mg -ma Normal=500N - 50*(6/5) Normal=440N

OpenStudy (anonymous):